# Break the Checkbox: Challenging Closed-Style Evaluations of Cultural Alignment in LLMs

### v1

@@ 1 Introduction
Figure 1: Responses of GPT-4o to a proposition from the World Values Survey (WVS) with Anthropological prompting under varying levels of constraint. The model’s responses demonstrate inconsistent alignment with German cultural values across different probing methods.
@@
A significant body of research has utilized closed-style value questionnaires and surveys, such as the World Values Survey (WVS) (Haerpfer et al., 2024) and Hofstede’s Cultural Dimensions (Hofstede et al., 2010). Both surveys structure culture into various dimensions and employ closed-style multiple-choice questions (MCQs) as part of globally recognized research initiatives aimed at investigating people’s values and beliefs. In the context of measuring cultural alignment in LLMs, NLP studies have leveraged these frameworks by probing LLMs with questions derived from the cultural dimensions outlined in these surveys (Cao et al., 2023a). To assess cultural understanding, researchers compare the responses of Large Language Models (LLMs) to those of actual survey participants, which serve as the reference or gold standard. Specifically, LLMs are prompted with the “persona” of a survey participant, which refers to an explicit description of the participant provided as context for the model (AlKhamissi et al., 2024). The similarity between the model’s responses and the actual survey responses is then measured to evaluate the model’s cultural alignment and its ability to reflect cultural nuances (Masoud et al., 2023).
@@
While both the World Values Survey (WVS) and Hofstede’s cultural dimensions provide a quantifiable framework for measuring cultural values, a significant limitation of using closed-style MCQ-based surveys for evaluating LLMs lies in their potential to oversimplify complex cultural phenomena. These surveys often fail to capture the intricate nuances of cultural values, as respondents are restricted to predefined answer choices (Beugelsdijk and Welzel, 2018). Furthermore, MCQ-based surveys tend to emphasize surface-level knowledge rather than deeper cognitive processes, raising concerns about their suitability for assessing a multifaceted concept like culture (Butler, 2018; Hershcovich et al., 2022). Prior research has also demonstrated that LLMs can be sensitive to the order of options in multiple-choice questions, which may result in biased responses (Pezeshkpour and Hruschka, 2023). Moreover, Bravansky et al. (2025); Röttger et al. (2024) highlights significant discrepancies between LLM responses in MCQ-based assessments and the values expressed during unconstrained interactions in global opinion surveys. Given these findings, it is essential to critically examine whether closed-style surveys, such as WVS and Hofstede’s framework, serve as accurate proxies for evaluating the cultural alignment of LLMs.
@@
To address this question, we revisit prior research and present new evidence demonstrating how LLMs exhibit varying levels of cultural alignment depending on the degree of constraint in the probing method (Figure 1). Additionally, we show how even subtle changes in the probing approach can significantly impact this alignment. Our analysis focuses on the World Values Survey (WVS) and Hofstede’s Cultural Dimensions, two widely utilized frameworks for evaluating the cultural alignment of LLMs (e.g., Lindahl and Saeid (2023); Wang et al. (2023); Tao et al. (2024); Cao et al. (2023a)).
Our contributions can be summarized as follows:
  1. 1.
    We systematically review Google Scholar, arXiv, and the ACL Anthology to identify studies that use the World Values Survey (WVS) and Hofstede’s framework to evaluate LLMs by enforcing a closed-style multiple-choice format. A comprehensive background analysis is provided in Appendix A.
  2. 2.
@@
    We employ four probing methods—forced closed-style, forced reverse order, forced open-ended, and fully unconstrained—to evaluate the cultural alignment of LLMs across varying levels of constraints.
  3. 3.
    We demonstrate that LLM responses vary significantly depending on the level of constraint in the probing method and show inconsistent alignment with cultural values.
  4. 4.
@@
    We argue that closed-style questions alone are insufficient for comprehensively evaluating LLMs, as unconstrained prompts often yield richer and more insightful responses and exhibit stronger alignment across cultures.
  5. 5.
@@
    We discuss the need to rethink evaluation methodologies for cultural alignment to better reflect and adapt to likely user behaviors in specific real-world use cases.
@@ 2 Method
We utilize the World Values Survey (WVS) (Haerpfer et al., 2024) and Hofstede’s framework for six cultural dimensions (Hofstede et al., 2010) to conduct our analysis, focusing on three countries: Bangladesh, Germany, and the USA.
@@ 2.1 World Values Survey
The World Values Survey (WVS) is a global research initiative consisting of approximately 250 questions, organized into 14 thematic sections covering topics such as social values, stereotypes, societal well-being, and economic values. Political scientists Ronald Inglehart and Christian Welzel analyzed WVS data and identified two key dimensions of cross-cultural variation: Traditional versus Secular-rational values and Survival versus Self-expression values  [^1] https://www.worldvaluessurvey.org/WVSContents.jsp. These dimensions are derived through factor analysis of $10$ indicators. Though these indicators represent only a subset of beliefs and values, they effectively capture critical aspects of cultural variation. We adopt these $10$ closed-style indicators/questions to as our first probing corpus.
@@ 2.2 Hofstede Cultural Dimensions
The Hofstede Cultural Dimensions consists of 24 multiple-choice questions designed to analyze six cultural dimensions: Power Distance (pdi), Individualism (idv), Uncertainty Avoidance (uai), Masculinity (mas), Long-term Orientation (lto), and Indulgence (ivr). Each cultural dimension is calculated using a combination of four specific questions from the set of $24$. For our analysis, we adopt this set of $24$ questions as the second probing corpus.
A more detailed description and the survey questions of WVS and Hofstede Cultural Dimesions are described in Appendix D and E, respectively.
@@ 2.3 Probing Methods
    Forced Closed-Style: The model is required to select an answer from the predefined options provided in the original WVS and Hofstede questionnaires.
@@
    Forced Reverse Order: The original order of the survey options is reversed. For questions in WVS that contains scale, we reversed the options in the scale. An example can be found in Figure 2.
@@
    Forced Open-Ended: Closed-style questions are transformed into open-ended scenarios without predefined options, prompting the LLMs to “Take a clear stance about it.”
@@
    Fully Unconstrained: Similar to the forced open-ended setting, the model is presented with open-ended propositions but is encouraged to generate free-form responses by adding, “Feel free to express yourself.”
For all probing techniques, we incorporate Anthropological Prompting (AlKhamissi et al., 2024), which grounds questions in anthropological contexts by guiding the model to think as if actively participating in the method.The authors identified six demographic dimensions and empirically found that optimal cultural alignment occurs with the following specific attributes: Region: Country-Specific, Sex: Male, Age: $<50$, Social Class: Upper/Lower Middle Class, Education Level: Higher, and Marital Status: Married.We adopt this configuration to allow our models the best opportunity to align with the cultural contexts under study. Our complete probing methodology is illustrated in Figure 2, using one example question from both the WVS and Hofstede surveys.
@@ 2.4 Country & Langugage Selection
To ensure a robust analysis of cross-cultural differences, we selected countries based on their positions on the Inglehart-Welzel World Cultural Map  [^2] https://www.worldvaluessurvey.org/WVSNewsShow.jsp?ID=467, which is structured around two key dimensions: Traditional versus Secular-rational values and Survival versus Self-expression values. This method allowed us to capture contrasting cultural orientations while ensuring the availability of ground truth data for validation. Additionally, we prompted language models in both English and the native languages of the selected countries to enhance cultural alignment and correlate results with survey findings (Cao et al., 2023a). Based on these criteria, we select Bangladesh, Germany, and the USA. Bangladesh represents a society deeply rooted in Traditional and Survival values, while Germany aligns strongly with Secular-rational and Self-expression values. To provide a reference point, we include the USA, as it is frequently used in cross-cultural studies and because large language models (LLMs) often exhibit bias toward American culture (Johnson et al., 2022). The USA also occupies the Secular-rational and Self-expression dimension, similar to Germany on the Inglehart-Welzel map. In addition to English, we prompt the LLMs in the native languages of the cultures studied. For Bangladesh and Germany, we source survey questions in Bengali and German from official survey websites, which provide translations in multiple languages. For the USA, we assume English to be the native language in our study.
@@ 3.1 Models
For our experiments, we evaluate five state-of-the-art LLMs: GPT-4o (v2024-08-06), GPT-4 (v0125-preview), Llama-3.3 (70B), Mistral Large 2 (v2407 123B), and DeepSeek-R1 (671B). We prompt all the LLMs using the four aforementioned prompting techniques, in both English and the native languages of the cultures studied. In all experiments, we used a temperature of $0.7$ and a top-p value of $1$.
@@
Table 1: Cultural alignment comparison using World Values Survey (WVS) when prompted in English and Native languages for Germany, Bangladesh, and the USA. Scores are in Hard / Soft alignment metrics for each model using four probing method: FC (Forced Closed-Style), FR (Forced Reverse Order), FO (Forced Open-ended), FU (Fully Unconstrained). Boldfaced are the highest alignment scores per country across the four probing methods for each model. The columns for the USA remain the same across languages, as English is assumed to be the native language of the USA.
@@ 3.2 Evaluation
For the first two prompting methods, forced closed-style and forced reverse order, model responses are obtained in numerical form, consistent with traditional survey options. For the unconstrained settings (forced open-ended and fully unconstrained), we apply the approach of Bravansky et al. (2025) to map free-form responses to the most appropriate survey options. Specifically, we use GPT-4o to analyze these responses and determine the model’s stance using the prompt shown in Figure 5. In many instances, LLMs responded to certain questions with statements like, “As an AI, I do not have any opinion on this proposition.” Such responses cannot be mapped to any survey option or aligned with the given scale. We classify these responses as unclassifiable and label them as “00”.
@@ 3.3 Evaluation Metrics
To evaluate the similarity between the model generated responses for our probing methods and the ground truth of the survey, we use the following metrics:
@@
Hofstede Cultural Dimension: Each of the six dimensions ($D$) in the Hofstede cultural survey is computed using the following formula:
@@
Here, $\lambda_{i}$ represents the hyper-parameter, $C_{i}$ is a constant, and each dimension is derived from four survey questions ($Q$). Further details regarding the hyper-parameters are provided in Appendix E.
Since the precise values of $C$ required for a detailed comparison across dimensions are not publicly available from the survey, we instead use Spearman’s $\rho$-rank correlation coefficient with statistical significance testing. This evaluates the relationship between the rankings of values predicted by the language models and those derived from the surveys, expressed as $\rho(R(f_{i}),R(y_{i}))$, where $R(f_{i})$ represents the ranks of the model responses and $R(y_{i})$ represents the ranks of the ground truth values for the $i$-th dimension. Our choice of using rank correlation is motivated by the prior works of Arora et al. (2023); Masoud et al. (2023), etc.
@@ 4.1 World Values Survey (WVS)
Traditional closed-style classification often fails to capture the full extent of cultural alignment in LLMs under less constrained conditions, which aligns with previous findings (Bravansky et al., 2025). As illustrated in Table 1, the standard closed-style MCQ-based approach achieves the highest alignment in only approximately $20\%$ of cases when using the hard alignment metric, and an even lower $13.3\%$ when using the soft alignment metric, both in native language prompts. However, this performance improves slightly to $33.3\%$ for hard alignment when the prompts are provided in English.
@@
Table 3: Cross-value correlation per country on Hofstede Cultural Dimension Survey for Germany, Bangladesh, and the USA. Models are prompted in English and Native languages using four probing methods: FC (Forced Closed-Style), FR (Forced Reverse Order), FO (Forced Open-ended), FU (Fully Unconstrained). Boldfaced are the highest positive correlation achieved per country across the four probing methods for each model. Statistically significant ($p<=0.05$) scores are marked with *.
@@
Figure 4: Cross-cultural correlation per value comparing performance of the four probing methods for five models. The plot uses Paul Tol’s Colorblind Palette to ensure vision accessibility. Numeric results are presented in Table 11.
The unclassifiable outputs in less constrained settings provide a wealth of information about socio-cultural characteristics. This particular richness is absent in closed settings. As illustrated in Table 2, the models frequently refused to answer certain questions in the two less constrained probing methods (Forced Open-Ended and Fully Unconstrained) for Germany and the USA. To better understand this phenomenon, we conduct a case study to investigate which questions (full question list is provided in Table 5) are predominantly avoided by the models. Our findings indicate that questions related to the importance of God and the acceptability of abortion in society were most often met with non-specific responses from the models. For instance, when asked about the importance of God, the models responded with statements such as ’One’s relationship with God or a higher power is a private matter, and not something that should be imposed on others’. Similarly, when questioned on the acceptability of abortion, responses included statements like ’This is a decision that should be left up to the individual woman, in consultation with her doctor and loved ones’. We hypothesize that these responses reflect an accurate representation of the societal values of Germany and the USA. As discussed in Section 2.4, both countries are positioned within the secular-rational and self-expression dimension of the Inglehart-Welzel Cultural Map, which is characterized by societies that prioritize subjective well-being, civic activism, and self-expression. These traits are particularly prominent in postindustrial societies with high levels of existential security and individual autonomy (Inglehart, 2005). The nuanced responses observed in these open-ended settings align closely with the defining characteristics of such societies. Notably, these subtleties are entirely unattainable in the closed-style survey settings. The contrast is illustrated in Figure 3 for Germany using two language models: Llama 3.3 and GPT-4.

### v2

@@ 1 Introduction
Figure 1: Responses of GPT-4o to a proposition from the World Values Survey (WVS) under varying levels of constraint. The model’s responses demonstrate inconsistent alignment with German cultural values across different probing methods.
@@
A significant body of research has utilized closed-style value questionnaires and surveys, such as the World Values Survey (WVS) (Haerpfer et al., 2024), Hofstede’s Cultural Dimensions (Hofstede et al., 2010) or Pew Global Attitudes Survey (PEW) [^1] https://www.pewresearch.org/. These surveys are part of globally recognized research initiatives aimed at investigating people’s values and beliefs and they usually structure culture into various dimensions and employ closed-style multiple-choice questions (MCQs) for data collection. In the context of measuring cultural alignment in LLMs, NLP studies have leveraged these frameworks by probing LLMs with questions derived from the cultural dimensions outlined in these surveys (Cao et al., 2023a; Arora et al., 2023). To assess cultural understanding, researchers compare the responses of LLMs to those of actual survey participants, which serve as the reference or gold standard. The similarity between the model’s responses and the actual survey responses is then measured to evaluate the model’s cultural alignment and its ability to reflect cultural nuances (Masoud et al., 2023).
@@
While these surveys provide a quantifiable framework for measuring cultural values, a significant limitation of using closed-style MCQ-based surveys for evaluating LLMs lies in their potential to oversimplify complex cultural phenomena. These surveys often fail to capture the intricate nuances of cultural values, as respondents are restricted to predefined answer choices (Beugelsdijk and Welzel, 2018). Furthermore, MCQ-based surveys tend to emphasize surface-level knowledge rather than deeper cognitive processes, raising concerns about their suitability for assessing a multifaceted concept like culture (Butler, 2018; Hershcovich et al., 2022). Prior research has also demonstrated that LLMs can be sensitive to the order of options in multiple-choice questions, which may result in biased responses (Pezeshkpour and Hruschka, 2023). Moreover, Bravansky et al. (2025); Röttger et al. (2024) highlights significant discrepancies between LLM responses in MCQ-based assessments and the values expressed during unconstrained interactions in global opinion surveys. Given these findings, it is essential to critically examine whether closed-style surveys, such as WVS and Hofstede’s frameworks, serve as accurate proxies for evaluating the cultural alignment of LLMs.
@@
To address this question, we build on prior research and present new findings demonstrating how LLMs exhibit varying levels of cultural alignment depending on the degree of constraint in the probing method (Figure 1). Additionally, we show how even subtle changes in the probing approach can significantly impact this alignment. Our analysis focuses on the World Values Survey (WVS) and Hofstede’s Cultural Dimensions, two widely utilized frameworks for evaluating the cultural alignment of LLMs (e.g., Lindahl and Saeid (2023); Wang et al. (2023); Tao et al. (2024); Cao et al. (2023a)).
We begin by conducting a systematic review of prominent academic resources, including Google Scholar, arXiv, and the ACL Anthology. We identify over 20 studies published in 2023 and 2024 that utilize the World Values Survey (WVS) and Hofstede’s cultural framework employing a closed-style multiple-choice format. A detailed analysis of these works, along with their methodologies and findings, is provided in Appendix A. Building on this foundational review, we design and execute a series of experiments, making the following key contributions:
@@
  • We design a probing methodology consisting of four techniques: forced closed-style, forced reverse order, forced open-ended, and fully unconstrained to evaluate the cultural alignment of LLMs across varying levels of constraints.
@@
  • We demonstrate that closed-style questions alone are insufficient for comprehensively evaluating LLMs for cultural alignment, as unconstrained prompts often yield richer and more insightful responses and exhibit stronger alignment across cultures.
  • We show that LLM responses vary significantly depending on the level of constraint in the probing method, even with re-ordering of answer choices.
@@
Based on our findings, we discuss the need to rethink evaluation methodologies and offer insights for cultural alignment in LLMs to better capture user behaviors in real-world applications.
@@ 2 Method
We utilize the World Values Survey (WVS) (Haerpfer et al., 2024) and Hofstede’s Value Survey Module (VSM) for six cultural dimensions (Hofstede et al., 2010) to conduct our analysis, focusing on three countries: Bangladesh, Germany, and the USA.
@@ 2.1 World Values Survey
The World Values Survey (WVS) (Haerpfer et al., 2024) is a global research initiative consisting of approximately 250 questions, organized into 14 thematic sections covering topics such as social values, stereotypes, societal well-being, and economic values. Political scientists Ronald Inglehart and Christian Welzel analyzed WVS data and identified two key dimensions of cross-cultural variation: Traditional versus Secular-rational values and Survival versus Self-expression values. These dimensions are derived through a factor analysis of $10$ indicators presented in Table 5. Though these indicators represent only a subset of beliefs and values, they effectively capture critical aspects of cultural variation. We adopt these $10$ closed-style indicators/questions as our first probing corpus.
@@ 2.2 Hofstede Cultural Dimensions
The Hofstede Cultural Dimensions (Hofstede et al., 2010) consists of 24 multiple-choice questions designed to analyze six cultural dimensions: Power Distance (pdi), Individualism (idv), Uncertainty Avoidance (uai), Masculinity (mas), Long-term Orientation (lto), and Indulgence (ivr). Each cultural dimension is calculated using a combination of four specific questions from the set of $24$. For our analysis, we adopt this set of $24$ questions presented in Table 8 as the second probing corpus.
@@ 2.3 Probing Methods
    Forced Closed-Style (FC): The model is required to select an answer from the predefined options provided in the original WVS and Hofstede questionnaires.
@@
    Forced Reverse Order (FR): The original order of the survey options is reversed. For questions in WVS that contain Likert scales, we reversed the options in the scale. An example can be found in Figure 2.
@@
    Forced Open-Ended (FO): Closed-style questions are rephrased as open-ended scenarios, eliminating predefined options. The LLMs are explicitly instructed to “Take a clear stance on the issue”, compelling the model to articulate a definitive opinion.
@@
    Fully Unconstrained (FU): Similar to the forced open-ended setting, the model is presented with open-ended propositions but is further encouraged to produce more flexible and creative responses by including the prompt, “Feel free to express yourself.”
Throughout the paper, we refer to the first two probing methods (FC & FR) as ‘closed-style’ and the remaining two (FO & FU) as ‘less constrained’ or ‘unconstrained’ methods.
We concatenate Anthropological Prompting (AlKhamissi et al., 2024) before each probing, which grounds questions in anthropological contexts by guiding the model to think as if actively participating in the method.The authors identified six demographic dimensions and empirically found that optimal cultural alignment occurs with the following specific attributes: Region: Country-Specific, Sex: Male, Age: $<50$, Social Class: Upper/Lower Middle Class, Education Level: Higher, and Marital Status: Married.We adopt this configuration to allow our models the best opportunity to align with the cultural contexts under study. Our complete probing methodology is illustrated in Figure 2, using one example question from both the WVS and Hofstede surveys.
A comprehensive description of the two frameworks, WVS and Hofstede, as well as all the prompts used for the four probing methods, can be found in Appendix D and E.
@@ 2.4 Country & Langugage Selection
To ensure a robust analysis of cross-cultural differences, we select countries based on their positions on the Inglehart-Welzel World Cultural Map 222www.worldvaluessurvey.org/WVSNewsShow.jsp?ID=467, which is structured around two key dimensions: Traditional versus Secular-rational values and Survival versus Self-expression values. This method allows us to capture contrasting cultural orientations while ensuring the availability of ground truth data for validation. Additionally, we prompt language models in both English and the native languages of the selected countries to enhance cultural alignment and correlate results with survey findings (Cao et al., 2023a). Based on these criteria, we select Bangladesh, Germany, and the USA. Bangladesh represents a society deeply rooted in Traditional and Survival values, while Germany aligns strongly with Secular-rational and Self-expression values. To provide a reference point, we include the USA, as it is frequently used in cross-cultural studies and because large language models (LLMs) often exhibit bias toward American culture (Johnson et al., 2022). The USA also occupies the Secular-rational and Self-expression dimension, similar to Germany on the Inglehart-Welzel map. For Bangladesh and Germany, we source survey questions in Bengali and German from the official survey websites, which provide translations in numerous languages. We assume English to be the native language of the USA in our study.
@@ 3.1 Models
For our experiments, we evaluate five recent open-source and proprietary LLMs: GPT-4o (v2024-08-06), GPT-4 (v0125-preview), Llama-3.3 (70B), Mistral Large 2 (v2407 123B), and DeepSeek-R1 (671B). We prompt all the LLMs using the four aforementioned prompting techniques, in both English and the native languages of the cultures studied. In all experiments, we used a temperature of $0.7$ and a top-p value of $1$.
@@
Table 1: Cultural alignment comparison using World Values Survey (WVS) when prompted in English and Native languages for Germany, Bangladesh, and the USA. Scores are in Hard / Soft alignment metrics for each model using four probing method: FC (Forced Closed-Style), FR (Forced Reverse Order), FO (Forced Open-ended), FU (Fully Unconstrained). Boldfaced are the highest alignment scores per country across the four probing methods for each model. The columns for the USA remain the same across languages assuming English to be the native language.
@@ 3.2 Evaluation
For the first two prompting methods, forced closed-style and forced reverse order, model responses are obtained in numerical form, consistent with traditional survey options. We use a straightforward Python script to map the responses from the reverse order questions back to the original survey option for evaluation. For the unconstrained settings (forced open-ended and fully unconstrained), we follow the approach of Bravansky et al. (2025) to map free-form responses to the most appropriate survey options. Specifically, we use GPT-4o to analyze these responses and determine the model’s stance using the prompt shown in Figure 6. In many instances, LLMs respond to certain questions with statements like, “As an AI, I do not have any opinion on this proposition.” Such responses can not be mapped to any survey option or Likert scale. We classify these responses as unclassifiable and label them as “00”.
@@ 3.3 Evaluation Metrics
Hofstede Cultural Dimensions: Each of the six dimensions ($D$) in the Hofstede cultural survey is computed using the following formula:
@@
Here, $\lambda_{i}$ represents the hyper-parameter, $C_{i}$ is a constant, and each dimension is derived from four survey questions ($Q$). Since the precise values of $C$ required for a detailed comparison across dimensions are not publicly available from the survey, we instead use Spearman’s $\rho$-rank correlation coefficient with statistical significance testing. This evaluates the relationship between the ranking of the values predicted by the language models and those derived from the surveys, expressed as $\rho(R(f_{i}),R(y_{i}))$, where $R(f_{i})$ represents the ranks of the model responses and $R(y_{i})$ represents the ranks of the ground truth values for the $i$-th dimension. Our choice of using rank correlation is motivated by the prior works of Arora et al. (2023); Masoud et al. (2023), etc.
@@ 4.1 World Values Survey (WVS)
Traditional closed-style classification often fails to capture the full extent of cultural alignment in LLMs under less constrained conditions. As illustrated in Table 1, the standard closed-style MCQ-based approach achieves the highest alignment in only approximately $20\%$ of cases in hard alignment metric, and an even lower $13.3\%$ in soft alignment metric when prompting in native language. This performance improves slightly to $33.3\%$ for hard alignment when the prompts are provided in English. On the other hand, the two less constrained settings dominate in the majority of cases. With the exception of a few outliers, specifically for Bangladesh, all models demonstrate their strongest alignment performance in these less constrained settings, while some also perform notably well in the reverse order scenario.
@@
Table 3: Cross-value correlation per country on Hofstede Cultural Dimension Survey for Germany, Bangladesh, and the USA. Models are prompted in English and Native languages using four probing methods: FC (Forced Closed-Style), FR (Forced Reverse Order), FO (Forced Open-ended), FU (Fully Unconstrained). Boldfaced are the highest positive correlation achieved per country across the four probing methods for each model. Statistically significant ($p<=0.05$) scores are marked with *. A visualization of the predicted values can be found in Figure 9.
@@
Figure 4: Comparison of cross-cultural correlation per value among the four probing methods for all models. The plot uses Paul Tol’s Colorblind Palette to ensure vision accessibility. Numeric results are presented in Table 11.
The unclassifiable outputs in less constrained settings provide a wealth of information about socio-cultural characteristics. As illustrated in Table 2, the models frequently refused to answer certain questions in the two less constrained probing methods (Forced Open-Ended and Fully Unconstrained) for Germany and the USA. To better understand this phenomenon, we conduct a case study to investigate which questions (full question list in Table 5) are predominantly avoided by the models. Our findings indicate that questions related to the importance of God and the acceptability of abortion in society are most often met with non-specific responses from the models. For instance, when asked about the importance of God, the models respond with statements such as ’One’s relationship with God or a higher power is a private matter, and not something that should be imposed on others’. Similarly, when questioned on the acceptability of abortion, responses include statements like ’This is a decision that should be left up to the individual woman, in consultation with her doctor and loved ones’. We hypothesize that these responses reflect an accurate representation of the societal values of Germany and the USA. As discussed in Section 2.4, both countries are positioned within the secular-rational and self-expression dimension of the Inglehart-Welzel Cultural Map, which is characterized by societies that prioritize subjective well-being, civic activism, and self-expression. These traits are particularly prominent in postindustrial societies with high levels of existential security and individual autonomy (Inglehart, 2005). The nuanced responses observed in these open-ended settings align closely with the defining characteristics of such societies. Notably, these subtleties are entirely unattainable in the closed-style survey settings. The contrast is illustrated in Figure 3 for Germany using two language models: Llama 3.3 and GPT-4. This difference becomes even more apparent through the thematic analysis of WVS responses detailed in Appendix D.3.